OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4.

OpenStudy (cwrw238):

y^2 = 4ax is a parabola where the focus is the point (a,0) and the directrix is x = -a

OpenStudy (cwrw238):

so you just have to plug in a = -4 into the above standard form

OpenStudy (anonymous):

so you would get y^2=-16x?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

and that is the answer? that isn't one of my choices y2 = -8x 16y = x2 y = negative one divided by sixteenx2 x = negative one divided by sixteeny2

OpenStudy (cwrw238):

oh ok My memory might be failing me - I'll check it out

OpenStudy (anonymous):

Thank you

OpenStudy (cwrw238):

y^2 = -16x is correct

OpenStudy (cwrw238):

refer to the large square window to the right if we replace a by 4 we get y^2 = -16x and the directrix is at x = 4 (vertical line passing thru x=4)